Aries:

A cake has circumference of 25 1/7 inches. What is the area of the cake? Use 22/7 to approximate π. Round to the nearest hundredth. Enter your answer in the box. ________in.2

jhonyy9:

formula of circumference ?

Aries:

What is the area of the cake?

chuckthedonut:

Do you remember what the formula of circumference is-?

Aries:

yes

Aries:

c=\[\pi\]d

chuckthedonut:

@aries wrote:
c=\[\pi\]d
Good, now we need to plug in our numbers.

Aries:

c= 25 1/7 \[\pi\]=22/7 d= ?

Renne:

To find the area of the cake, we need to first find its radius. To do that, we can use the formula for the circumference of a circle: C = 2πr, where C is the circumference and r is the radius. Plugging in the given values, we get: 25 1/7 = 2 x 22/7 x r Simplifying, we get: r = (25 1/7) / (2 x 22/7) r = 3 8/21 Now that we have the radius, we can use the formula for the area of a circle: A = πr^2 Plugging in the given values, we get: A = 22/7 x (3 8/21)^2 A ≈ 29.14 Rounding to the nearest hundredth, we get: A ≈ 29.14 in.2 (Tried to explain as best I could)

Aries:

Okay, I get it. Thank you!

Renne:

No problem